OpenStudy (anonymous):

Patrick drew the incorrect graph shown below to represent the circle (x – 1)2 + (y + 3)2 = 4 Which statement explains why Patrick’s graph is incorrect? Answers A. The radius of the circle should be 2 instead of 4. B. The center of the circle should be (1, 9) instead of (1, -3). C. The center of the circle should be (-1, 3) instead of (1, -3). D. The radius of the circle should be 16 instead of 4.

OpenStudy (anonymous):

@PaxPolaris can you help me?

OpenStudy (paxpolaris):

a circle with center \((h,k)\) and radius \(r\) will have the formula: \[\Large \left( x-h \right)^2+\left( y-k \right)^2=r^2\]

OpenStudy (anonymous):

Okay

OpenStudy (paxpolaris):

so in your formula: what's the center supposed to be? and what's the radius is supposed to be?

OpenStudy (anonymous):

1, -3 r=4?

OpenStudy (paxpolaris):

center is (1,-3)... yes but \(\large r \ne 4\) \[\Large \color{green}{r^2=4}\]

OpenStudy (anonymous):

So r = 2?

OpenStudy (anonymous):

Or 16?

OpenStudy (paxpolaris):

2... you got it.

OpenStudy (anonymous):

Great, thanks a million!